OpenStudy (anonymous):

solve the inequality 0 ≤ 2 − 5(x + 1) < 27

OpenStudy (jdoe0001):

well, so, how would you solve 2 − 5(x + 1) = 27 ? well, you add this and multiply by that to the TWO sides well, in this case you do the same, but with THREE sides so, whenever you add/multiply to one, you do it to all THREE

OpenStudy (jdoe0001):

there's only one thing you'd need to bear in mind on that, whenever you use a NEGATIVE multiplier or divider or exponential like -1 you have to "flip" the inequality sign

OpenStudy (jdoe0001):

harobi7256 confused still?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok $$ \begin{matrix} 0 \le& 2 − 5(x + 1) &\lt 27\\ 0 \le & 2 − 5x -5 & \lt 27\\ 0 \le & − 5x -3 & \lt 27\\ \text{add 3 to isolate the "x"}\\ 3 \le& − 5x &\lt 30\\ \end{matrix} $$

OpenStudy (jdoe0001):

now you want to solve by "x", so you just need to get rid of the pesky -5 in the middle

OpenStudy (anonymous):

ok so be \[\frac{ -3 }{ 5 } \le x < -6\]

OpenStudy (jdoe0001):

yes, you divide by -5, or multiply by -(1/5) HOWEVER as I said above, whenever you use a negative multiplier or exponential, you have to "flip" the inequality sign

OpenStudy (anonymous):

okay thank you!

OpenStudy (jdoe0001):

$$ -\cfrac{3}{5}\ge x \gt 6 $$

OpenStudy (jdoe0001):

-6 I mean hehe $$ -\cfrac{3}{5}\ge x \gt -6 $$